# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v2

@@ 4.1.2 Topics of conversation with companion chatbots
Analysis of both structured responses and free-text data revealed five major themes in how users engage with companion chatbots (see Figure 7).
Casual exchange, including casual conversations (26.28%) and entertainment (21.66%), represents the most common form of interaction. Participants used the chatbot to review past events (12.60%), such as talking about their day: “I often chat with my AI about random things like my day, and it feels like I’m having a normal conversation without the pressure of talking to a real person.”
Emotional disclosure emerged as a notable theme, encompassing mental health discussions and personal problems. This is reflected in the quantitative findings of personal issues and mental health (14.17%) and interpersonal issues and drama (10.96%) as notable topics. One participant responded, “I can talk about my problems, and it’s like having a private conversation with no fear of being criticized.”
Though less prevalent in the free responses, the theme of creative development manifested through topics of writing, roleplay, and creative ideation, with users leveraging chatbots as collaborators: “I typically ask for ideas for my marketing art! I do a lot of graphic design for a lot of different people and sometime run low on creative ideas.”
@@
A rare theme emerging from the free responses was intimate exchange, involving romantic and sexual interactions: “[Some]times I just want to feel pleasure and the interaction is very sexual. Other times I want to feel love and connection and it’s more romantic.”
We note that the themes identified in free responses sometimes diverge from the structured response options. This diverse range of conversation topics highlights the versatility of companion chatbots in meeting users’ needs for entertainment, emotional support, and intellectual stimulation.
@@ 4.2.1 Correlations in psychological, social, and companion chatbot usage characteristics
For our prespecified analyses of usage vs. loneliness, we found a small but significant correlation between chatbot session length and loneliness ($r=0.101,p=0.042$), but not between loneliness and usage frequency ($r=0.053,p=0.287$), as shown in Figure 9.
Several other notable correlations emerged in our exploratory analysis using Benjamini-Hochberg corrected p-values. Chatbot usage (both length and frequency) showed moderate positive correlations with state empathy towards chatbots ($r>0.3,p<0.001$) and various forms of attraction towards chatbots ($r>0.18,p<0.001$). Loneliness demonstrated significant correlations with personality traits of extraversion, ($r=-0.188,p<0.001$), conscientiousness ($r=-0.335,p<0.001$), agreeableness ($r=-0.333,p<0.001$), and neuroticism ($r=0.526,p<0.001$), aligning with Buecker et al.’s [Buecker et al.(2020)] meta-analysis. Additionally, problematic use of chatbots (GPIUS2) showed moderate correlations with both session length ($r=0.303,p<0.001$) and frequency ($r=0.337,p<0.001$). See Figure 8 for a correlation matrix of our continuous variables.
@@
Figure 8: Correlation matrix of psychological, social, and usage characteristics. Variables are sorted into general categories, and names have been altered for readability. The original names can be seen in Figure 18 of Appendix A.
Figure 9: Plots of loneliness (ULS-8) and problematic usage (GPIUS2) vs. usage (frequency and session length). Top left: loneliness vs. chatbot usage frequency. Top right: loneliness vs. chatbot session length. Bottom left: problematic use vs. chatbot usage frequency. Bottom right: problematic use vs. chatbot session length..
Figure 10: The final model from our exploratory multiple regression, depicting the relationships between chatbot session length, loneliness, and related factors.
@@ 4.2.2 Multiple Regression Model
[…]
A parallel multiple mediation analysis (Figure 12) revealed that problematic internet use (GPIUS2) significantly mediated the relationship between session length and loneliness ($\beta=0.0183$, 95% CI $[0.0027,0.0403]$), while social network size (LSNS) did not show significant mediation ($\beta=0.0007$, 95% CI $[-0.0304,0.0340]$). The total indirect effect was not statistically significant ($\beta=0.0191$, 95% CI $[-0.0149,0.0563]$).
Figure 13: Model summary of an ordinary least squares regression for the effect of chatbot session length on loneliness (ULS-8), using standardized variables. Model Formula: Loneliness (ULS-8) $\sim$ Session Length Score + Social Attraction towards Chatbot + Problematic Use (GPIUS2) + Socialization (LSNS) + Social Attraction towards Close Person + Neuroticism + Agreeableness + Sex + Age + Session Length Score $\times$ Social Attraction towards Close Person
@@ 4.3.1 Overall characteristics
Participants ranged in age from 18 to 73, primarily 30-35. Most (93.6%) were from the United States, white/Caucasian (66.3%), and employed full-time (61.6%), with a broad distribution of household income. For relationship status, most were single (42.1%) or married (32.4%). 59.7% of participants identified as men and 36.6% identified as women. Additional demographic characteristics are in Figure 14.
Most participants utilize Snapchat MyAI (31.6%), Character.ai (30.4%), or Replika (22.6%) and tend to spend 15 and 30 minutes in each session with a companion chatbot (42.6%), with 35.4% reported sessions shorter than 15 minutes. Very few users spend extended periods chatting, with only 1.28% having sessions exceeding 4 hours. Most participants have been using their companion chatbots for a few months (43.9%) to around a year (35.2%), with only 7.7% of participants having used their chatbots for multiple years and 13.2% for around a month or less. More details on usage can be seen in Figure 4 and Figure 5.
@@ 4.3.2 Clusters of users of companion chatbots
Our K-means cluster analysis revealed seven distinct user profiles among companion chatbot users, each exhibiting unique behaviors, motivations, and psychological characteristics. The resulting clusters are visualized in the heatmap and dendrogram presented in Figure 15. The demographics of each cluster compared to overall demographics can be seen in Figure 14, with percentages found in Figure 20 of Appendix B.
Validation of the cluster solution using Kruskal-Wallis tests revealed significant differences among clusters in terms of session length, loneliness (ULS-8), problematic use of chatbots (GPIUS2), and frequency of use (all $p<0.001$). The average silhouette score ($-0.0413$) suggested some overlap between clusters, which is not uncommon in complex psychological data. A MANOVA test confirmed significant multivariate differences among the clusters when considering all variables simultaneously (Wilks’ $\lambda=0.2420,F(24,1375.71)=28.77,p<0.001$). All four MANOVA test statistics (Wilks’ lambda, Pillai’s trace, Hotelling-Lawley trace, and Roy’s greatest root) were significant ($p<0.001$), providing robust evidence for cluster differences across multiple dimensions. Box’s M-test indicated significant heterogeneity of covariance matrices ($\chi^{2}=181.96$, $p<0.001$), which does not invalidate the clustering but suggests the need for cautious interpretation. This heterogeneity may affect the reliability of post-hoc comparisons and could indicate that the relationship between variables differs across clusters. While these results provide strong support for the distinctiveness of the identified clusters, they should still be considered exploratory, providing a foundation for more targeted investigations in future research.
Figure 15: Right: Cluster means of standardized variables of chatbot usage (frequency and session length), loneliness, and problematic use of chatbots for our seven user clusters. Left: a dendrogram depicting the degree of similarity between each cluster.
@@
  • 0: Disengaged Light Users (7.67%): Minimal chatbot engagement with average loneliness levels. Show neutral attitudes toward AI technology and average social support levels. Primary usage is brief (<15 min) casual conversations, with main motivations being curiosity to start use and practical utility for continued use. Average age is 32; for gender, 65% are men and 35% are women.
@@
Detailed heatmaps including all numerical scales can be seen in Figure 16. A version with the original values of the scales can be found in Figure 19 along with detailed cluster descriptions in Appendix B. The patterns of themes for each cluster’s motivations for use and topics of discussion can be seen in Figure 6 and Figure 7.
Post-hoc analysis using Dunn’s test with Bonferroni correction revealed distinct patterns among clusters across our key variables (Figure 17). For session length, Fulfilled Dependent Users (Cluster 4) showed significantly longer sessions than all other clusters (p < .001) except Well-Adjusted Moderate Users. Disengaged Light Users (Cluster 0) had significantly shorter sessions than most other clusters (p < .001).
@@
Figure 17: Significant pairwise differences between clusters on key variables based on Dunn’s test with Bonferroni correction. Cell background color indicates significance of differences (darker green = presence of greater significance in differences), with text color indicating direction (red = greater than, blue = less than). For example, >2*** indicates significantly greater values than Cluster 2 at p < .001. Significance levels: *: p < .05, **: p < .01, ***: p < .001.
@@ 5 Discussion
Our study provides a detailed examination of the relationship between companion chatbot usage and loneliness, offering some insights into the complex interplay of psychological, social, and technological factors that shape this relationship.
@@ 5.1 Understanding the Relationship Between Companion Chatbot Usage and Loneliness
To answer RQ2, “What is the relationship between companion chatbot usage and loneliness, and what factors mediate or moderate this relationship?” we look to our regression analysis. While session length showed no significant direct relationship with loneliness, our regression analysis revealed a more nuanced picture through mediating and moderating factors. Problematic use emerged as a significant mediator, suggesting that increased usage may create more opportunities for problematic behaviors to develop. This aligns with existing research connecting problematic internet use as both a cause and effect of loneliness [Kim et al.(2009)], indicating a similar pattern may exist for companion chatbot use.
The negative association between usage and social network size warrants attention, though causality remains unclear. Users might reduce human interactions as they spend more time with chatbots, or they might turn to chatbots to fill an existing social void. The positive relationship between usage and social attraction towards chatbots is logically consistent – more interaction time allows deeper attachments to form. However, whether this attraction leads to loneliness through unfulfilled relational needs or loneliness enhances attraction to chatbots is uncertain.
Among confounding variables, neuroticism showed a strong positive association with loneliness, while agreeableness demonstrated a strong negative association, aligning with previous research on personality traits and loneliness [Buecker et al.(2020)], reinforcing the importance of controlling for these factors in studies of technology use and social well-being.
An interesting finding was the positive interaction term between session length and social attraction towards a close person. This suggests a nuanced relationship between chatbot usage and loneliness for socially attracted individuals. This could be interpreted as a mismatch between social needs and chosen social interactions, where time spent with chatbots may be perceived as time not spent with valued human connections. We can also consider a reversed causality: perhaps a very lonely user becomes more socially attracted to people as they spend more time with their chatbot, or non-lonely users could be satisfied with their chatbots and subsequently become less socially attracted to people.
@@ 5.2 Understanding Clusters of Chatbot Users
We shall first discuss the overall characteristics of our participants, which we explored for RQ1, “What motivates the use of companion chatbots, and what do people use them for?” Our analysis revealed that despite companion chatbots’ marketing as social tools, most participants primarily used them for technological exploration and entertainment rather than friendship or intimate relationships. This usage pattern may parallel other solitary activities like gaming or reading, which can affect socialization and lead to overindulgence. However, our cluster analysis revealed distinct user profiles demonstrating more complex relationships with these technologies. This brings us to RQ3, “Can we identify and characterize distinct user profiles based on usage patterns, loneliness levels, problematic use, and other psychological and social factors?”
@@
The most striking contrast emerges between Socially Fulfilled Dependent Users (Cluster 4) and Lonely Moderate Users (Cluster 3). Both exhibit high problematic use, yet their outcomes differ dramatically. Socially Fulfilled Dependent Users maintain high well-being despite (or perhaps because of) their extensive chatbot engagement. Their high empathy and attraction towards chatbots, combined with positive reported effects on real-world social interactions, suggests that intensive chatbot use can coexist with healthy social functioning. These findings parallel work by Snodgrass et al. showing that more intensively involved gamers, those exhibiting what appears to be “addictive” play, experience the greatest psychosocial benefits from their gaming – including reduced loneliness [Snodgrass et al.(2018)]. However, unlike gaming where more intense involvement tends to predict better social outcomes, we found that Lonely Moderate Users with similar usage patterns experienced markedly different results. For chatbots, the nature of engagement may matter more than intensity alone. This challenges concerns about the negative effects of AI replacing human relationships [Pataranutaporn et al.(2021), Turkle(2011)] and merits further investigation into how these users maintain this balance.
Conversely, Lonely Moderate Users demonstrate how similar usage patterns can lead to markedly different outcomes. Their high loneliness, low socialization, and preference for chatbots over humans raise concerns about whether chatbot use exacerbates social withdrawal or merely reflects existing social difficulties. Their pattern of emotional disclosure to chatbots suggests these tools may serve as their primary outlet for emotional expression. We suggest further investigation of whether chatbots have a positive or negative effect on these individuals.
@@
Socially Challenged Frequent Users (Cluster 6) and Lonely Light Users (Cluster 5) are another interesting pair to consider. Socially Challenged Frequent Users are most similar to Lonely Moderate Users (3), though they have average problematic use and considerably below-average measures of trust and empathy, even for those they feel close to. These individuals likely need intervention beyond what a companion chatbot can currently provide, yet companion chatbots may be the only entities they are willing to engage with. Both Socially Challenged Frequent Users and Lonely Light Users seek out chatbots for companionship, and it is largely only in these two groups that participants discussed having or seeking a romantic or sexual relationship with their chatbot. The participants in these two clusters most strongly reflect the Media Equation, in that their responses to chatbots may be similar to their social responses to people [Reeves and Nass(1998)]. Of concern is whether or not Lonely Light Users might transition to Socially Challenged Frequent Users over time, losing trust and empathy with people as they engage more with chatbots. Considering that loneliness is a risk factor for problematic use [Kim et al.(2009)], understanding how these users’ characteristics can change over time, and how chatbot systems can detect these at-risk individuals and intervene, may be valuable for future research.
@@
Well-Adjusted Moderate Users (Cluster 1), the largest group, demonstrate a potentially optimal engagement pattern. With moderate usage and lower emotional connection to AI, they maintain strong human connections while leveraging chatbots primarily as tools for emotional processing rather than social companions. This suggests that chatbots might effectively reduce loneliness without necessarily functioning as relationship substitutes. Many disclose their emotional issues to chatbots, valuing the safe space for expression they provide, with a few considering the chatbot as part of their lifestyle. Chatbots might effectively reduce loneliness without necessarily functioning as relationship substitutes. Further investigation into causality and mechanisms of psychosocial changes may prove fruitful for understanding how companion chatbots can be designed to reduce loneliness, besides the typical strategy of pushing for greater emotional intimacy with the user. For consideration, recent work has shown that social support acts as a significant mediator between psychological problems and problematic gaming [Malak et al.(2023)], which may help explain how Cluster 4, with poor perceived social support, exhibits high loneliness – as opposed to Cluster 1, with high perceived social support and low loneliness – when both have moderate chatbot usage.
@@
The AI-Wary Light Users (Cluster 2) provide an important counterpoint, showing that emotional distance from chatbots can coexist with low loneliness. These users show low emotional attachment to chatbots and low opinions of AI in general. This group uses chatbots mostly for practical purposes and recreation, perhaps without seeking or developing emotional connections. Many traits associated with loneliness – neuroticism, agreeableness, socialization, and self-esteem) are fairly average, but their loneliness is the second lowest of all clusters. Considering this in comparison to other clusters suggests that emotional investment in AI companions may not be necessary or even optimal for all users.
These diverse profiles highlight that the impact of companion chatbot use on loneliness and social well-being varies significantly based on individual characteristics and usage patterns. This heterogeneity suggests the need for personalized approaches in both chatbot design and potential intervention strategies.
5.3 Implications and Future Directions
Our findings extend beyond simple social displacement theories [Kraut et al.(1998)]suggesting a framework where companion chatbots serve different social functions based on user characteristics and usage patterns. This aligns with and extends the Parasocial Contact Hypothesis [Schiappa et al.(2005)], which suggests that media characters can provide meaningful social experiences. Chatbot interactions may similarly provide parasocial benefits, but with important distinctions based on user traits and usage patterns.
The mediating role of problematic use in the relationship between usage and loneliness contributes to theories of technology addiction and compensatory internet use [Kardefelt-Winther(2014)]. Excessive chatbot use may reflect attempts to compensate for social deficits, similar to problematic social media use, but with unique characteristics due to the interactive nature of companion chatbots. This extends existing frameworks for understanding problematic technology use to include AI companionship specifically.

### v3

@@ 4.1.2 Topics of conversation with companion chatbots
Analysis revealed five major themes in chatbot engagement (see Figure˜10 in the Appendix). Our automated analysis confirmed these themes while providing prevalence estimates across all 404 participants. The convergence between manual and automated approaches strengthened confidence in our thematic findings.
Casual exchange, including casual conversations (26.28%) and entertainment (21.66%), represents the most common interaction: “I often chat with my AI about random things like my day, and it feels like I’m having a normal conversation without the pressure.”
Emotional disclosure also emerged as notable, encompassing mental health discussions (14.17%) and interpersonal issues (10.96%): “I can talk about my problems, and it’s like having a private conversation with no fear of being criticized.”
@@
The theme of creative development manifested through topics of writing, roleplay, and creative ideation, with users leveraging chatbots as collaborators: “I typically ask for ideas for my marketing art! I do a lot of graphic design for a lot of different people and sometime run low on creative ideas.” A rare theme was intimate exchange, involving romantic and sexual interactions.
This diverse range of conversation topics highlights the versatility of companion chatbots in meeting users’ needs for entertainment, emotional support, and intellectual stimulation.
Figure 3: A: The final model from our exploratory multiple regression, depicting the relationships between chatbot session length, loneliness, and related factors. B: Model summary of an ordinary least squares regression for the effect of chatbot session length on loneliness (ULS-8), using standardized variables.
@@
Figure 4: Plots of loneliness (ULS-8) and problematic usage (GPIUS2) vs. usage (frequency and session length). Top left: loneliness vs. chatbot usage frequency. Top right: loneliness vs. chatbot session length. Bottom left: problematic use vs. chatbot usage frequency. Bottom right: problematic use vs. chatbot session length.
@@ 4.2.1 Correlations in psychological, social, and companion chatbot usage characteristics
For our prespecified analyses of usage vs. loneliness, we found a small but significant correlation between chatbot session length and loneliness ($r=0.101,p=0.042$), but not between loneliness and usage frequency ($r=0.053,p=0.287$), as shown in Figure˜4.
Several other notable correlations emerged in our exploratory analysis using Benjamini-Hochberg corrected p-values. Chatbot usage (both length and frequency) showed moderate positive correlations with state empathy towards chatbots ($r>0.3,p<0.001$) and various forms of attraction towards chatbots ($r>0.18,p<0.001$). Loneliness demonstrated significant correlations with personality traits of extraversion, (r=0.188,p<0.001r=-0.188,p<0.001), conscientiousness (r=0.335,p<0.001r=-0.335,p<0.001), agreeableness (r=0.333,p<0.001r=-0.333,p<0.001), and neuroticism ($r=0.526,p<0.001$), aligning with Buecker et al.’s [80] meta-analysis. Additionally, problematic use of chatbots (GPIUS2) showed moderate correlations with both session length ($r=0.303,p<0.001$) and frequency ($r=0.337,p<0.001$). See Figure˜1 for a correlation matrix of our continuous variables.
@@ 4.2.2 Multiple Regression Model
We conducted an ordinary least squares (OLS) regression analysis to examine the relationship between session length and loneliness while controlling for other factors. The model explained a significant proportion of the variance in loneliness ($R^{2}=0.496$, adjusted $R^{2}=0.482$, $F(11,392)=35.09$, $p<.001$) (Figure˜3 and Figure˜3).
While session length alone did not significantly predict loneliness ($\beta=0.060$, $p=0.126$), neuroticism ($\beta=0.358$, $p<.001$) and problematic chatbot use ($\beta=0.166$, $p<.001$) showed positive associations with loneliness. Social network size ($\beta=-0.316$, $p<.001$) and agreeableness ($\beta=-0.147$, $p<.001$) showed negative associations. We found a significant interaction effect (see Figure˜7 in the Appendix) between session length and social attraction to a close person ($\beta=0.054$, $p=0.049$). This interaction suggests that for individuals with higher social attraction to close others, longer chatbot sessions were associated with slightly increased loneliness, while this relationship was absent or reversed for those with lower social attraction to close others.
A parallel multiple mediation analysis (Figure˜8 in the Appendix) revealed that problematic internet use (GPIUS2) significantly mediated the relationship between session length and loneliness ($\beta=0.0183$, 95% CI $[0.0027,0.0403]$), while social network size (LSNS) did not show significant mediation ($\beta=0.0007$, 95% CI [0.0304,0.0340][-0.0304,0.0340]). The total indirect effect was not statistically significant ($\beta=0.0191$, 95% CI [0.0149,0.0563][-0.0149,0.0563]).
@@ 4.3.1 Overall characteristics
Participants ranged in age from 18 to 73, primarily 30-35. Most (93.6%) were from the United States, white/Caucasian (66.3%), and employed full-time (61.6%), with a broad distribution of household income. For relationship status, most were single (42.1%) or married (32.4%). 59.7% of participants identified as men and 36.6% identified as women. Additional demographic characteristics are in Figure˜11.
Most participants utilize Snapchat MyAI (31.6%), Character.ai (30.4%), or Replika (22.6%) and tend to spend 15 and 30 minutes in each session with a companion chatbot (42.6%), with 35.4% reported sessions shorter than 15 minutes. Very few users spend extended periods chatting, with only 1.28% having sessions exceeding 4 hours. Most participants have been using their companion chatbots for a few months (43.9%) to around a year (35.2%), with only 7.7% of participants having used their chatbots for multiple years and 13.2% for around a month or less. More details on usage can be seen in Figure˜2.
@@ 4.3.2 Clusters of users of companion chatbots
Our K-means cluster analysis revealed seven distinct user profiles among companion chatbot users, each exhibiting unique behaviors, motivations, and psychological characteristics. The resulting clusters are visualized in the heatmap and dendrogram presented in Figure˜5. The demographics of each cluster, compared to the overall demographics, are shown in Figure˜11, with percentages provided in Figure˜11 of the Appendix.
Validation of the cluster solution using Kruskal-Wallis tests revealed significant differences among clusters in terms of session length, loneliness (ULS-8), problematic use of chatbots (GPIUS2), and frequency of use (all $p<0.001$). The average silhouette score (0.0413-0.0413) suggested some overlap between clusters, which is not uncommon in complex psychological data. A MANOVA test confirmed significant multivariate differences among the clusters when considering all variables simultaneously (Wilks’ $\lambda=0.2420,F(24,1375.71)=28.77,p<0.001$). All four MANOVA test statistics (Wilks’ lambda, Pillai’s trace, Hotelling-Lawley trace, and Roy’s greatest root) were significant ($p<0.001$), providing robust evidence for cluster differences across multiple dimensions. Box’s M-test indicated significant heterogeneity of covariance matrices ($\chi^{2}=181.96$, $p<0.001$), which does not invalidate the clustering but suggests the need for cautious interpretation. This heterogeneity may affect the reliability of post-hoc comparisons and suggests that the relationship between variables varies across clusters. While these results provide strong support for the distinctiveness of the identified clusters, they should still be considered exploratory, providing a foundation for more targeted investigations in future research.
Figure 5: Left: a dendrogram depicting the degree of similarity between each cluster. Right: Cluster means of standardized variables of chatbot usage (frequency and session length), loneliness, and problematic use of chatbots for our seven user clusters. The cluster order corresponds to that of the dendrogram.
@@
  • 0: Disengaged Light Users (7.67%): Minimal chatbot engagement with average loneliness levels. Show neutral attitudes toward AI technology and average social support levels. Primary usage is brief (less than 15 min) casual conversations, with main motivations being curiosity to start use and practical utility for continued use. Average age is 32; for gender, 65% are men and 35% are women.
@@
Detailed heatmaps including all numerical scales can be seen in Figure˜12. A version with the original values of the scales can be found in Figure˜13 along with detailed cluster descriptions in the Appendix. The patterns of themes for each cluster’s motivations for use and topics of discussion can be seen in Figure˜10 in the Appendix.
Post-hoc analysis using Dunn’s test with Bonferroni correction revealed distinct patterns among clusters across our key variables (Figure˜9). For session length, Fulfilled Dependent Users (Cluster 4) showed significantly longer sessions than all other clusters ($p<.001$) except Well-Adjusted Moderate Users. Disengaged Light Users (Cluster 0) had significantly shorter sessions than most other clusters ($p<.001$).
@@ 5 Discussion
Our study examines the relationship between companion chatbot usage and loneliness, revealing complex interactions among psychological, social, and technological factors.
@@ 5.1 Understanding the Relationship Between Companion Chatbot Usage and Loneliness
For RQ2, our regression analysis showed no significant direct relationship between session length and loneliness, but revealed important mediating and moderating factors. Problematic use emerged as a significant mediator, suggesting increased usage creates more opportunities for problematic behaviors, aligning with research connecting problematic internet use with loneliness [81].
Usage is negatively associated with social network size, though causality remains unclear—users might reduce human interactions as they spend more time with chatbots or turn to chatbots to fill existing social voids. The positive relationship between usage and social attraction towards chatbots is logically consistent, as interaction time allows deeper attachments to form.
Among confounding variables, neuroticism showed strong positive association with loneliness, while agreeableness demonstrated strong negative association, aligning with previous research [80] and reinforcing the importance of controlling for these factors.
The positive interaction between session length and social attraction towards close people suggests a nuanced relationship, possibly indicating a mismatch between social needs and chosen interactions, where time with chatbots may be perceived as time not spent with valued human connections.
@@ 5.2 Understanding Clusters of Chatbot Users
For RQ1, most participants used companion chatbots primarily for technological exploration and entertainment rather than friendship or intimate relationships. Our cluster analysis for RQ3 revealed distinct user profiles demonstrating complex relationships with these technologies.
@@
The most striking contrast emerged between Socially Fulfilled Dependent Users (Cluster 4) and Lonely Moderate Users (Cluster 3). Both exhibit high problematic use, yet with dramatically different outcomes. Socially Fulfilled Dependent Users maintain high well-being despite extensive chatbot engagement, paralleling findings that intensively involved gamers experience psychosocial benefits from their gaming [82]. This challenges concerns about AI replacing human relationships [44, 8]. Conversely, Lonely Moderate Users with similar usage patterns experience high loneliness, low socialization, and preference for chatbots over humans.
@@
Socially Challenged Frequent Users (Cluster 6) and Lonely Light Users (Cluster 5) present another interesting comparison. Both seek chatbots for companionship, with participants discussing romantic or sexual relationships with their chatbots. These clusters most strongly reflect the Media Equation [83]. Of concern is whether Lonely Light Users might transition to Socially Challenged Frequent Users over time, losing trust and empathy with people. Understanding how these users’ characteristics change over time may be valuable for future research.
@@
Well-Adjusted Moderate Users (Cluster 1), the largest group, demonstrate a potentially optimal engagement pattern, maintaining strong human connections while using chatbots primarily as tools for emotional processing. Many disclose emotional issues to chatbots, valuing the safe space they provide. Recent work showing social support as a significant mediator between psychological problems and problematic gaming [84] may help explain how Cluster 3, with poor perceived social support, exhibits high loneliness compared to Cluster 1.
@@
The AI-Wary Light Users (Cluster 2) provide an important counterpoint, showing that emotional distance from chatbots can coexist with low loneliness. These users show low emotional attachment to chatbots and use them mostly for practical purposes, suggesting emotional investment in companion chatbots may not be necessary or optimal for all users.
5.3 Ethical Considerations & Design Implications
Our findings raise critical questions about responsible companion chatbot development. The identification of vulnerable user populations who may develop problematic usage patterns highlights the need for proactive ethical safeguards. The contrast between users who benefit from AI companionship and those who may be harmed underscores that one-size-fits-all approaches may be ethically insufficient.
Protecting Vulnerable Users. Clusters 3, 5, and 6 represent particularly vulnerable populations who often turn to AI companions for emotional support and companionship, yet may be most susceptible to negative outcomes. Current platforms typically lack mechanisms to identify such users or provide appropriate interventions. Our findings suggest the need for risk assessment tools that identify users showing patterns associated with problematic use, adaptive interfaces that promote healthy usage patterns for vulnerable users, and intervention mechanisms such as usage limits, mental health resource suggestions, or encouragement of human social contact.
Developer Responsibility & Regulation. Unlike traditional software tools, AI companions can explicitly encourage emotional attachment and dependency, creating ethical obligations similar to those in healthcare or counseling contexts. This includes implementing safeguards for user design principles focused on encouraging and facilitating human social connections. The largely unregulated nature of AI companion platforms, combined with their potential for both benefit and harm, suggests the need for evidence-based policy consideration. Our findings could inform regulatory approaches that balance innovation with user protection, particularly for vulnerable populations.